# Robust Multi-View Learning via Representation Fusion of Sample-Level Attention and Alignment of Simulated Perturbation

### v1

@@ 8 More Experimental Results
Table 7: Performance comparison of unsupervised multi-view clustering on multi-view datasets (mean±$\pm$±std)
@@
Table 8: Performance comparison of unsupervised multi-view clustering on multi-view datasets (mean±$\pm$±std)
@@
Table 9: Performance comparison on noise-label multi-view classification (mean±$\pm$±std)
@@
Regarding hyper-parameter $\lambda$, we consider noise-label multi-view classification and cross-modal hashing retrieval tasks, where $\mathcal{L}_{\rm RML}$ is treated as a regularization term weighted by $\lambda$. The parameter analysis with the noise label rate of $50\%$ is shown in Figure 7, where we observe stable classification performance within the range of $[10^{1},10^{2},10^{3}]$. For the noise-label multi-view classification task, $\lambda$ is set to $10^{3}$ to emphasis $\mathcal{L}_{\rm RML}$ in joint optimization when the noise label rates are large (e.g., $30\%$, $50\%$, $70\%$). When the noise label rates are small (e.g., $0\%$, $10\%$), $\lambda$ is set to $10^{0}$ for recommended settings. For the cross-modal hashing retrieval tasks, stable performance is observed within the range of $[10^{-3},10^{-2},10^{-1}]$ as shown in Figure 8. On cross-modal retrieval datasets MIRFLICKR-25K and NUS-WIDE, we kept $\lambda$ unchanged in our comparison experiments (i.e., $\lambda=10^{-1}$).

### v2

@@ 8 More Experimental Results
Table 7: Performance comparison of unsupervised multi-view clustering on multi-view datasets (mean $\pm$ std)
@@
Table 8: Performance comparison of unsupervised multi-view clustering on multi-view datasets (mean $\pm$ std)
@@
Table 9: Performance comparison on noise-label multi-view classification (mean $\pm$ std)
@@
Regarding hyper-parameter $\lambda$, we consider noise-label multi-view classification and cross-modal hashing retrieval tasks, where $\mathcal{L}_{\rm RML}$ is treated as a regularization term weighted by $\lambda$. The parameter analysis with the noise label rate of $50\%$ is shown in Figure 7, where we observe stable classification performance within the range of $[10^{1},10^{2},10^{3}]$. For the noise-label multi-view classification task, $\lambda$ is set to $10^{3}$ to emphasize $\mathcal{L}_{\rm RML}$ in joint optimization when the noise label rates are large (e.g., $30\%$, $50\%$, $70\%$). When the noise label rates are small (e.g., $0\%$, $10\%$), $\lambda$ is set to $10^{0}$ for recommended settings. For the cross-modal hashing retrieval tasks, stable performance is observed within the range of $[10^{-3},10^{-2},10^{-1}]$ as shown in Figure 8. On cross-modal retrieval datasets MIRFLICKR-25K and NUS-WIDE, we kept $\lambda$ unchanged in our comparison experiments (i.e., $\lambda=10^{-1}$).